OpenStudy (anonymous):

Solving: 3x^2 = -12x -15 Do you subtract 3x^2 from both sides? ..I did that and the whole problem was negative, but when someone else did it, it was positive.

OpenStudy (anonymous):

It is better to addn12x +15 to the left side. then use quad formula or try to factor

Directrix (directrix):

3x^2 = -12x -15 3x^2 +12x +15 = 0 3(x^2 + 4x + 5) = 0 Now, what?

OpenStudy (anonymous):

So, do you always have to make it positive? Or does it depend on the problem?

Directrix (directrix):

Positive has nothing to do with this. It just turned out that way. In solving a quadratic by factoring, the technique is to gather all terms on one side of the equation with 0 on the other. Your task now is to factor or get the roots of x^2 + 4x + 5. I think you'll need to use the Quadratic Formula.

OpenStudy (anonymous):

But why couldn't it work out the way I had done it in the first place, where I subtracted 3x^2 from both sides.

Directrix (directrix):

0mgl0la Let's see what happens. 3x^2 = -12x -15 0 = -3x^2 -12x -15 0 = -3( x^2 + 4x + 5) Divide both sides by -3. 0 = x^2 + 4x + 5